# Effective Game-Theoretic Motion Planning via Nested Search

### v1

@@ 7 Conclusion and future work
In the longer run, we plan to develop extensions of the approach to the closed-loop setting , and study principled tie-breaking and fairness-/risk-aware objectives, as well as coordination mechanisms when several NEs exist.
Appendix 0.A Additional Figures
Table 2: Notation for continuous and graph–based multi–robot planning.
@@ Appendix 0.A Additional Figures
$x^{i}(t)$ $x(t)=(x^{1}(t),\dots,x^{m}(t))$ Robot $i$ (joint) state at time $t\in[0,T]$
$u^{i}(t)$ $u(t)=(u^{1}(t),\dots,u^{m}(t))$ Robot $i$ (joint) control at time $t\in[0,T]$
@@
$\pi^{i}_{x_{0}^{i},u^{i}}$ $\pi_{x_{0},u}=\times_{i=1}^{m}\pi^{i}_{x_{0}^{i},u^{i}}$ Robot $i$ (joint) continuous trajectory
$0,T]\mapsto\mathcal{X}^{i}$ $\bigl([0,T]\to\mathcal{X}\bigr)$
$\Pi^{i}$ $\Pi=\times_{i=1}^{m}\Pi^{i}_{n}$ All robot $i$ (joint) feasible continuous
@@
$c^{i}(\pi^{i}(t),\pi^{-i}(t))$ Stage cost of robot $i$ given others’
trajectories
$c(\pi(t))$ Stage cost over joint trajectory
$J^{i}(\pi^{i},\pi^{-i})$ Cost of robot $i$ given others’ trajectories
@@
$\bigl(e=(x,y)\bigr)$ with control
$u^{i}_{e^{i}}\in\mathcal{U}^{i}_{\Delta t}$ $\bigl(u_{e}\in\mathcal{U}_{\Delta t}\bigr)$
$\pi^{i}_{e^{i}}$ $\pi_{e}=(\pi^{1}_{e^{1}},\ldots,\pi^{m}_{e^{m}})$ Robot $i$ (joint) trajectory along edge $e^{i}$ $\bigl(e\bigr)$
$\pi^{i}_{n}{=}(x^{i}_{e^{i}_{1}},\dots,x^{i}_{e^{i}_{n}})$ $\pi_{n}{=}(x_{e_{1}},\dots,x_{e_{n}})$ Robot $i$ (joint) $n$-step trajectory
@@
$\Pi^{i}_{n}$ $\Pi_{n}=\times_{i=1}^{m}\Pi^{i}_{n}$ Robot $i$ (joint) feasible $n$-step trajectories
@@
(a) Local NE Solver
(b) Global NE Solver
Figure 8: Second-order bicycle model for car dynamics. Car dimensions $l,w,L$ in black; State variables $(x,y,\theta,v,\delta)$ in red.
Figure 9: Kinodynamic track graph built from racetrack centerline with lateral offsets. Node, edge styling as in grid-graph.

### v2

@@ 7 Conclusion and future work
In the longer run, we plan to extend the approach to the closed-loop setting and study principled tie-breaking and fairness- and risk-aware objectives, as well as coordination mechanisms when several NEs exist.
Appendix 0.A Implementation Speedups
As mentioned in Sec. 4, the pseudocode is presented so that the fundamentals of GTNS and isNashEquilibrium are as clear as possible. In practice, we obtain significant efficiency improvements by leveraging several implementation techniques.
The most consequential (inspired by works such as [dellin2016unifying, mandalika2019generalized, cohen2015planning]) is lazy validation of collision and equilibrium checks: every neighbour of $x$ visited in line 1 is pushed to the OPEN list without immediate validation; only when a neighbour is popped from OPEN (i.e., it currently has the lowest cost) do we verify that it is in fact a valid node. This defers the most expensive computations and spends resources only on promising candidates.
The implementation also caches edge-collision results (Boolean flags indicating whether an edge $e\in E$ is in collision) to allow fast lookup before invoking an expensive check, precomputes heuristics, and performs early feasibility checks for the query. Additional optimizations include careful memory management and the use of pointers to stable storage. Finally, we note that isNashEquilibrium is trivially parallelizable, as the per-robot A* runs with respect to fixed trajectories of all other robots are independent.
Appendix 0.B Additional Experiments & Figures
This appendix includes supplementary figures and experiments that exceed the space limitations of the main manuscript. Specifically, we examine the effects of NE mismatch and its resulting adverse outcomes. Additionally, we present an experiment that isolates the influence of $\lambda_{\mathrm{prox}}$ in a semi-collaborative setting. Finally, a summary of the notation used in the paper is provided.
0.B.1 Effect of NE mismatch
In the majority of tested cases, when agents executed in an open-loop fashion trajectories belonging to different NEs, collisions between robots were encountered. This underscores the need for a coordinated, team-wide equilibrium in behavior-aware motion planning. To address the concern that open-loop per-agent NEs are “obviously brittle,” we also evaluated an MPC variant: at every step each agent re-solved its own NE instance, where $J=J_{i}$, i.e., the global cost strictly optimizes the solution of the specific robot executing its trajectory from this NE. At every step, the robot resolved the problem towards the goal using this objective while observing the updated locations of the other robots, executed its step, and repeated the process. Such an approach is hoped to be more resilient as it could provide more opportunities for the robots to avoid colliding with each other. Unfortunately, even under this closed-loop regime, the lack of shared equilibrium selection caused systematic issues and in almost all scenarios (e.g., those in Figs. 4 and 1) a collision. In rare “best-case” scenarios, when interaction geometry was simple, (e.g., on the track in Fig. 5 with large start distance) mismatched individual NEs yielded similar solutions to those returned by the global planner.
[…]
Figure 5: Following distance. Semi-collaborative setting, such as two cars on the same team, with social welfare cost $J=\sum_{i}J^{i}$ and $\lambda_{\mathrm{prox}}\in[0,1]$. The cars are attempting to cross the finish line, with the lowest possible summed cost. The addition of a proximity penalty to the cost, affects the following distance that Robot 1 (blue) maintains from Robot 2 (orange): A small $\lambda_{\mathrm{prox}}$ value yields close drafting, and larger values increase the time-gap as Robot 1 deliberately backs off from Robot 2, visible at the first corner.
(a) Local NE Solver
(b) Global NE Solver
Figure 7: Kinodynamic track graph built from racetrack centerline with lateral offsets. At each longitudinal sample we define a wayline (normal to the centerline), set the heading from the centerline tangent, and discretize $(v,\delta)$. From each wayline $\ell$ we attempt connections from all nodes on $\ell$ to all nodes on any wayline $\ell^{\prime}$ with $|\ell^{\prime}-\ell|\leq k$. Each discrete state $(x,y,\theta,v,\delta)$ is drawn with two arrows anchored at $(x,y)$: a blue heading arrow pointing in direction $\theta$ (fixed short length), and a red steering arrow pointing in direction $\theta+\delta$ with length $\propto v$. Edge styling as in grid-graph.
Figure 8: Second-order bicycle model for car dynamics. Car dimensions $l,w,L$ in black; State variables $(x,y,\theta,v,\delta)$ in red. In our experiments, we used a scale-model car of length $l=0.70\,\mathrm{m}$, width $w=0.20\,\mathrm{m}$, and wheelbase $L=0.50\,\mathrm{m}$.
Figure 9: Generic maps: with (bottom) and without (top) obstacles.
Table 2: Notation for continuous and graph-based multi-robot planning.
@@ 0.B.1 Effect of NE mismatch
$x^{i}(t)$ $x(t){=}(x^{1}(t),...,x^{m}(t))$ Robot $i$ (joint) state at time $t\in[0,T]$
$u^{i}(t)$ $u(t){=}(u^{1}(t),...,u^{m}(t))$ Robot $i$ (joint) control at time $t\in[0,T]$
@@
$\pi^{i}_{x_{0}^{i},u^{i}}$ $\pi_{x_{0},u}{=}(\pi^{1}_{x_{0}^{1},u^{1}},...,\pi^{m}_{x_{0}^{m},u^{m}})$ Robot $i$ (joint) continuous trajectory
$\pi^{i}:[0,T]\mapsto\mathcal{X}^{i}$ $\bigl(\pi:[0,T]\mapsto\mathcal{X}\bigr)$
$\Pi^{i}$ $\Pi=\times_{i=1}^{m}\Pi^{i}$ All robot $i$ (joint) feasible continuous
@@
$c^{i}(\pi^{i}(t),\pi^{-i}(t))$ Stage cost of robot $i$ given its and
others’ trajectories
$c(\pi(t))$ Joint stage cost over joint trajectory
$J^{i}(\pi^{i},\pi^{-i})$ Cost of robot $i$ given its and
others’ trajectories
@@
$\bigl(e=(x,y)\bigr)$ with control $u^{i}_{e^{i}}\in\mathcal{U}^{i}_{\Delta t}$
$\bigl(u_{e}\in\mathcal{U}_{\Delta t}\bigr)$
$\pi^{i}_{e^{i}}$ $\pi_{e}{=}(\pi^{1}_{e^{1}},...,\pi^{m}_{e^{m}})$ Robot $i$ (joint) trajectory on edge $e^{i}\ \bigl(e\bigr)$
$\pi^{i}_{n}{=}(x^{i}_{e^{i}_{1}},...,x^{i}_{e^{i}_{n}})$ $\pi_{n}{=}(x_{e_{1}},...,x_{e_{n}})$ Robot $i$ (joint) $n$-step trajectory
@@
$\pi^{i}_{x^{i}}$ $\pi_{x}{=}(\pi^{1}_{x^{1}},...,\pi^{m}_{x^{m}})$ Robot $i$ (joint) trajectory to state $x^{i}\ \bigl(x\bigr)$
$\pi^{i}_{x^{i}\to y^{i}}$ $\pi_{x\to y}$ Robot $i$ (joint) local trajectory from
state $x^{i}$ to $y^{i}\ \bigl(x$ to $y\bigr)$
$\Pi^{i}_{n}$ $\Pi_{n}=\times_{i=1}^{m}\Pi^{i}_{n}$ Robot $i$ (joint) feasible $n$-step
trajectories
$P^{i}(x^{i})$ $P(x)$ Predecessor map of node $x^{i}\ \bigl(x\bigr)$